OpenStudy (anonymous):

Can someone help me solve this using elimination? 2x + 3y = 4 4x + 6y = 9 First I would turn the first equation into 4x+6y=8 or would it be 8x + 3y = 4?

OpenStudy (anonymous):

Do you multiply the 4 from the second equation with the entire second equation or only with the x variable?

OpenStudy (anonymous):

4x + 6y = 8 is a correct way to use elimination on your first equation.

OpenStudy (anonymous):

Could you explain the reasoning behind that? I kind of just guessed... @kmcllb

OpenStudy (jhannybean):

begin by working to eliminate the x first. 2x + 3y = 4 4x + 6y = 9 _____________ In order to do that we need to multiply the first equation by -2 -2(2x +3y = 4) 4x +6y =9

OpenStudy (anonymous):

oh, so I don't have to multiply it by the 4 as long as one of the variables has the same value in both equations?

OpenStudy (anonymous):

Well, in elimination you are trying to either "cross" out your X value or Y value, The way you guessed, you are eliminating your Y value. In order to elimination your values need to be the same number. So all you did was multiply that by x2 to get your new equation; now just eliminate!

OpenStudy (jhannybean):

YEah, you're trying to "eliminate" one variable to solve for the other

OpenStudy (jhannybean):

therefore when you multiply-2 tot he first equation and then simplify it, you'll be adding the two equations together and thus the x value will be eliminated.

OpenStudy (anonymous):

@Jhannybean how come you used a negative 2 when multiplying the equation?

OpenStudy (jhannybean):

because I am trying to "eliminate" the x to solve for the y variable.

OpenStudy (anonymous):

Because to eliminate you NEED a NEGATIVE AND POSITIVE Y or X

OpenStudy (jhannybean):

One min brb

OpenStudy (anonymous):

so it would be -4x - 6y = -8 4x + 6y = 9

OpenStudy (jhannybean):

Looks like there is no solution to the system of equations, unless we did something wrong. Hmm.

OpenStudy (anonymous):

Yes, now choose which one you would like to eliminate, X or Y.

OpenStudy (anonymous):

so then I subtract right?

OpenStudy (anonymous):

0x = 1, the answer is simply, "No solution" and yes that's what math does... they tell you there's always a answer but over time some of them are never solvable.

OpenStudy (jhannybean):

If you subtract them both x and y will become 0, therefore no solution.

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

thank you for the help, I was also wondering if you could help guide me through a word problem?

OpenStudy (jhannybean):

I could try. Close this question and open a new one :)

OpenStudy (anonymous):

Yes, I will guide you, but you will have to conclude the answer among yourself

OpenStudy (anonymous):

Roast beef has 25 g o protein and 11g of calcium per serving. A serving of mashed potatoes has 2g of protein and 25g of calcium How many servings of each are needed to supply exactly 29 g of protein and 61 g of calcium?

OpenStudy (anonymous):

Sir, you add them... Beef Stats + Potatoes stats = 29G and 61 Calcium....

OpenStudy (anonymous):

The answer is: 1 Serving of Roast Beef, 2 Servings of Mashed Potatoes.

OpenStudy (anonymous):

Could you help me write that out as the equations though, that's where I don't understand it

OpenStudy (anonymous):

25G + 11C + 2(2G + 25C) = 29G + 61C

OpenStudy (anonymous):

oh, thank you so much, sorry for the trouble

OpenStudy (anonymous):

It's ok, I excel in math so for me this is easy, but I understand the struggle for you. If you live in America, Happy Thanksgiving! :)